# ZipGAN: Super-Resolution-based Generative Adversarial Network Framework for Data Compression of Direct Numerical Simulations

### v2

@@ 3.2 Progressive transfer learning approach
Figure 5: Number of snapshots (#$\mathrm{\#\,snapshots}$) required (bars) and training time ($\mathrm{T_{training}}$) (solid line) needed to achieve the accuracy levels reported in Fig. 4 (left), with and without the use of transfer learning (TL), on the Re90 dataset.
One challenge is the computational expense of saving the original DNS snapshots, and this challenge is compounded by the lengthy training time for SR-GAN models. To mitigate these issues, transfer learning (TL) (see Sec. 2.1), is applied for $\mathrm{n_{\Delta}>2}$. TL significantly reduces the number of required snapshots and, as a result, decreases the overall training time, particularly at higher upsampling factors. By leveraging previously acquired knowledge, TL decreases training time to achieve ¯$\bar{\mathcal{E}}({|\nabla\mathbf{u}|})<0.05$ by approximately $50\%$ with every doubling of $\mathrm{n_{\Delta}}$. TL is clearly effective in enhancing training efficiency and reducing computational demands.
@@ 3.3 Enhancing temporal resolution
Compression method Orig. DWT DWT DWT   ZipGAN DWT   ZipGAN DWT   ZipGAN DWT   ZipGAN
[…]
@@
Figure 6: Left: Correlation of the quality score, $\mathrm{Q_{s}}$, with both the NSME of the velocity gradients, ¯$\bar{\mathcal{E}}({|\nabla\mathbf{u}|})$ (top frame), and the SSIM (bottom frame) obtained from the $\phi_{\mathrm{dec}}$ using ZipGAN across different compression ratios, $\theta$. Right: Qualitative illustration of the SR-GAN-based approach for enhancing temporal resolution. Original DNS fields ($\mathrm{\phi_{orig}}$, purple squares) are saved at intervals, filtered, and downsampled (purple arrows) to create in-sample encoded fields ($\mathrm{\phi_{enc}}$, purple circles) for SR-GAN training. Out-of-sample fields are only saved as encoded fields ($\mathrm{\phi_{enc}}$, red circles) and are decoded by the SR-GAN to enhance temporal resolution. The pretrained discriminator provides a quality score ($\mathrm{Q_{s}}$) which is used to assess both in-sample (purple asterisks) and out-of-sample (red asterisks) decoded fields.
@@ 3.4 Decoding time
Compression method DWT DWT DWT   ZipGAN DWT   ZipGAN DWT   ZipGAN DWT   ZipGAN
$T_{\Gamma}$ [s] – Re90 0.29 0.29 0.31  0.013 0.32  0.062 0.34  0.28 0.34  1.63
$T_{\Gamma}$ [s] – Re350 5.84 5.87 5.88  0.038 5.89  0.35 5.91  1.76 5.92  13.74
Table 4: Comparison of the decoding time, $\mathrm{T_{\Gamma}}$, using both the DWT and ZipGAN compression methods applied to the Re90 and Re350 datasets for different compression ratios. The DWT employs 48 CPUs and ZipGAN uses a single GPU.
@@ 4 Conclusion
We investigate the performance of ZipGAN, a SR-GAN-based compression and decompression method, for the storage and transfer of DNS HIT datasets and compare it with the performance of the DWT method. The ZipGAN architecture, with its generator acting as a decoding operator, accurately reconstructs high-resolution fields from their encoded, low-resolution counterparts. Once trained, the ZipGANś generator can decode fields solely from the encoded input, i.e., without requiring further access to the original DNS data.
@@
In summary, ZipGAN offers a highly efficient and scalable solution for compressing large DNS datasets of turbulent flows without introducing large decoding errors at high compression ratios. It is particularly well suited for applications requiring the preservation of fine-scale features and sufficient temporal resolution such as three-dimensional turbulent flow datasets.
In order to enhance the reproducibility of this study, provide clarity on technical aspects, and facilitate faster development, access to our GIT repository [^2] https://git.rwth-aachen.de/nista.ludovico/DataCompression_ZipGAN.git will be granted upon request.

### v3

@@ 3.2 Progressive transfer learning approach
Figure 5: Number of snapshots ($\mathrm{\#\,snapshots}$) required (bars) and training time ($\mathrm{T_{training}}$) (solid line) needed to achieve the accuracy levels reported in Fig. 4 (left), with and without the use of transfer learning (TL), on the Re90 dataset.
One challenge is the computational expense of saving the original DNS snapshots, and this challenge is compounded by the lengthy training time for SR-GAN models. To mitigate these issues, transfer learning (TL) (see Sec. 2.1) is applied for $\mathrm{n_{\Delta}>2}$. TL significantly reduces the number of required snapshots and, as a result, decreases the overall training time, particularly at higher upsampling factors. By leveraging previously acquired knowledge, TL decreases training time to achieve $\bar{\mathcal{E}}({|\nabla\mathbf{u}|})<0.05$ by approximately $50\%$ with every doubling of $\mathrm{n_{\Delta}}$. TL is clearly effective in enhancing training efficiency and reducing computational demands.
@@ 3.3 Enhancing temporal resolution
Compression
method
Orig. DWT DWT
DWT   ZipGAN
DWT   ZipGAN
DWT   ZipGAN
DWT   ZipGAN
$\mathrm{Q_{s}}$ $\mathrm{[\%]}$ – Re90
99.9 99.7 99.2
36.6    98.8
9.3      96.7
1.6      95.6
0.8      88.8
@@
Figure 6: Left: Correlation of the quality score, $\mathrm{Q_{s}}$, with both the NSME of the velocity gradients, $\bar{\mathcal{E}}({|\nabla\mathbf{u}|})$ (top frame), and the SSIM (bottom frame) obtained from the $\phi_{\mathrm{dec}}$ using ZipGAN across different compression ratios, $\theta$. Right: Qualitative illustration of the SR-GAN-based approach for enhancing temporal resolution. Original DNS fields ($\mathrm{\phi_{orig}}$, purple squares) are saved at intervals, filtered, and downsampled (purple arrows) to create in-sample encoded fields ($\mathrm{\phi_{enc}}$, purple circles) for SR-GAN training. Out-of-sample fields are only saved as encoded fields ($\mathrm{\phi_{enc}}$, red circles) and are decoded by the SR-GAN to enhance temporal resolution. The pretrained discriminator provides a quality score ($\mathrm{Q_{s}}$) which is used to assess both in-sample (purple asterisks) and out-of-sample (red asterisks) decoded fields.
@@ 3.4 Decoding time
Compression
method
DWT DWT
DWT   ZipGAN
DWT   ZipGAN
DWT   ZipGAN
DWT   ZipGAN
$T_{\Gamma}$ [s] – Re90
0.29 0.29
0.31  0.013
0.32  0.062
0.34  0.28
0.34  1.63
$T_{\Gamma}$ [s] – Re210
5.84 5.87
5.88  0.038
5.89  0.35
5.91  1.76
5.92  13.74
Table 4: Comparison of the decoding time, $\mathrm{T_{\Gamma}}$, using both the DWT and ZipGAN compression methods applied to the Re90 and Re210 datasets for different compression ratios. The DWT employs 48 CPUs and ZipGAN uses a single GPU.
@@ 4 Conclusion
We investigate the performance of ZipGAN, a SR-GAN-based compression and decompression method, for the storage and transfer of DNS HIT datasets and compare it with the performance of the DWT method. The ZipGAN architecture, with its generator acting as a decoding operator, accurately reconstructs high-resolution fields from their encoded, low-resolution counterparts. Once trained, the ZipGAN’s generator can decode fields solely from the encoded input, i.e., without requiring further access to the original DNS data.
@@
In summary, ZipGAN offers a highly efficient and scalable solution for compressing large DNS datasets of turbulent flows without introducing large decoding errors at high compression ratios. It is particularly well-suited for applications requiring the preservation of fine-scale features and sufficient temporal resolution, such as three-dimensional turbulent flow datasets.
In order to enhance the reproducibility of this study, provide clarity on technical aspects, and facilitate faster development, access to our GIT repository will be granted upon request.